Show that the complete integral of z = px +qy - 2p - 3q represent all possible planes through the points (2, 3, 0)
1
Expert's answer
2020-03-12T12:39:05-0400

Given that

"z=px+qy-2p-3q\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (1)"

Its complete integral is


"z=ax+by-2a-3b\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (2)"


"a, b" being arbitrary constants.

Differentiating "(2)" partially with respect to "x"


"{\partial z \over \partial x}=a=>p=a"

Differentiating "(2)" partially with respect to "y"


"{\partial z \over \partial y}=b=>q=b"

Substituting in "(2)" we get


"z=px+qy-2p-3q"

The equation "(2)" is a linear equation in "x,y,z." Hence it represents planes for various values of 𝑎 and 𝑏.  Substitute "x=2, y=3, z=0"


"2a+3b-2a-3b=0=>0=0"

The coordinates of the point (2, 3, 0) satisfy the equation "(2)." Hence the complete integral "(2)" of "(1)" represents all possible planes passing through the point (2, 3, 0).